OpenStudy (anonymous):

can I use 180 instead of pi in this formula?

OpenStudy (anonymous):

\[A = \pi^{o} - \arcsin \frac{ b \ sinA }{ \alpha }\]

OpenStudy (anonymous):

well pie is 3. a bunch of numbers are you converting from radians to deggres or degrees to radians?

OpenStudy (anonymous):

sorry, \[\pi\] not \[\pi ^{o}\]

OpenStudy (anonymous):

I know,I know pi is 180 but i'm not sure that I can use 180 instead of pi.

OpenStudy (dape):

\(\pi≠180\) but \(\pi=180°\). So if you make sure you measure everything in degrees the formula would still work.

OpenStudy (dape):

That also means using the degree versions of sin and arcsin.

OpenStudy (anonymous):

oh ok , Thank you very much :)